Graph the slope 2x+6=3y
Answer by algebrahouse.com(1659)

Slope-intercept form is y = mx + b
m is the slope
b is the y-intercept

2x + 6 = 3y
y = (2/3)x + 2
slope = 2/3
y-intercept = 2
.

To ask a question, visit: www.algebrahouse.com